Previous | Calendar |
September |
1999 |
Next | |||||||||||||||||||||||
---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday | |||||||||||||||||||||
29JD241 |
30JD242 |
31JD243 |
1JD244 |
2JD245 |
3JD246Site Visit (AT): Details span tank had dropped 200 psi (leak in new valve?) turned off the 1-m T/RH sensor (H2O was getting into the connector) turned off fast o3 pump Start of AmeriFlux intercomparison (with Bob Evans) |
4JD247at 7:00, data running fine |
|||||||||||||||||||||
5JD248Site Visit (AT): Details from 11:00-11:15, manual calibration of EC/LI-6262 system from 11:15-11:20, swapped prof/LI-6251 dessicant at 12:15, 2-m T/RH running again at 13:25, o3 turned off at 13:44, manual cal of EC/LI-6262 system |
6JD249 |
7JD250 |
8JD251Site Visit (AT): Details from 14:00-15:00, put new SS inlet on CO2 line; cleaned krypton hygrometer at 16:00, new filters on fast.o3 line; turned on fast1.o3 (not both) from 16:00-17:00, moved ATI sonic to 13m End of AmeriFlux intercomparison (with Bob Evans) |
9JD252 |
10JD253Site Visit (AT): Details from 10:30-11:18 EC system down (leak at inlet to IRGA), fixed! manual cal of LI-6262 |
11JD254
|
12JD255
|
13JD256
|
14JD257
|
15JD258Site Visit (AT): Details at 15:00, changed o3 target for fast1.o3 from 15:15-15:30, prof system down
|
16JD259
|
17JD260
|
18JD261
|
19JD262
|
20JD263Yom Kippur
|
21JD264Site Visit (AT): Details from 10:00-15:00, re-arranged radiation sensors and data logger connections Blanken 23X --> start using CNR3.CSI (CNR-1, PAR, NR.LITE, REBS Q*7.1) light21x --> met21x.csi using MET.CSI (prop-vane winds, RH1-3, baro pressure) at 14:00, changed prep.config at 15:30, autocals OFF
|
22JD265Site Visit (AT): Details covars didn't run correctly (error in covar.config from yesterday, fixed, re-ran covars) from 11:30-16:00, added dessicant trap to profile system (seemed to be problematic, see AT logbook) Fall Equinox
|
23JD266
|
24JD267Site Visit (AT): Details at 10:00, swapped dessicant trap attached to UHP N2 cylinder from 10:00-13:00, moved profiler inlets, L1=0.5m; L2=1m; L3=2m; L4=5m; L5=10.5m; L6=21.5m; changed inlet filters on L1-L4, still need to change L5 + L6
|
25JD268
|
26JD269
|
27JD270
|
28JD271
|
29JD272
|
30JD273Site Visit (AT): Details from 11:30-11:45, changed desiccant on profiler
|
1JD274
|
2JD275 |